OpenStudy (anonymous):

I tried everything and I still can't figure out the answer. Can anyone tell me what it is please? Evaluate the expression below when x = -2 x^2-8/x+6 A) -3 B) -1 C) 1/2 D) 3/2

OpenStudy (igreen):

Oops, my bad. It said x = -2.

OpenStudy (igreen):

Input -2 for all values of x: \(\dfrac{x^2-8}{x+6}\) \(\dfrac{-2^2-8}{-2+6}\) Simplify Exponent: \(\dfrac{4-8}{-2+6}\) Subtract & Add: \(\dfrac{-4}{4}\) Simplify: \(-1\)

OpenStudy (igreen):

Do you get how I did that? @JulianoRossiel

OpenStudy (anonymous):

Omg I get it now, it's so obvious.. You get a medal for sure (: And more medals to come